For the following data, find the correlation co-efficient. How much 𝒙 depends on π’š?

Determine and sketch the corresponding regression line. Graphically find the value of 𝒙

when π’š = πŸπŸ“.


  • 𝒙: 5 12 18 23 27 30 26 22
  • π’š: 18 16 13 11 9 7 10 13




1
Expert's answer
2021-10-04T16:36:57-0400

Solution:


X Values

βˆ‘ = 163

Mean = 20.375

βˆ‘(X - Mx)2Β = SSxΒ = 489.875


Y Values

βˆ‘ = 97

Mean = 12.125

βˆ‘(Y - My)2Β = SSyΒ = 92.875


X and Y Combined

NΒ = 8

βˆ‘(X - Mx)(Y - My) = -208.375


R Calculation

r = βˆ‘((X - My)(Y - Mx)) / "\sqrt{(SS_x)(SS_y)}"

r = -208.375 / "\sqrt{(489.875)(92.875)}" = -0.9769

This is a strong negative correlation, which means that high X variable scores go with low Y variable scores (and vice versa).

Sketch of regression line:



From this graph, when y=15, x=13.5